OpenStudy (anonymous):

Solve for y, then find the value of y when given x = –2. 6x = 7 – 4y

OpenStudy (jamesj):

When x = -2 6x = -12 So 12 = 7 - 4y hence ... can you take it from here?

OpenStudy (anonymous):

idk i was doing this problem and i got that far and i messed it up i think. it came out like 10-y and i was like what? im awful at math. ugh. :(

OpenStudy (anonymous):

To solve for y, subtract 7 from both sides and then divide both sides by -4. \[y = (-6x + 7) \div 4\] then just plug in x value when x = -2: \[y = [-6(-2) +7] \div 4 \] y(-2) = 4.75

OpenStudy (anonymous):

oops i forgot a negative?

OpenStudy (anonymous):

wait but the options are -12, 7 fourths, 19 fourths, and 24.

OpenStudy (anonymous):

y = 19/4 = 4.75 I didn't know it was multiple choice, I just put it into decimal form